k110 9 +5=14 -4 n-3=-11 8 17-3 C) 10-8=-24 3 nd 9、 the (L) The quotient of a number and-7, decreased by 2, is 10. Find the number.
Answer #1

Answer: (J) k=99, (V) m=-36, (H) n=-64, (O) w=100, (T) x=-48, (E) a=160 and (L) Number is -84.

Explanation:

(J)

1 10

1110+1

\frac{k}{9}=11

9\times \frac{k}{9}=9\times 11

k=99

(V)

\frac{m}{-4}+5=14

\frac{m}{-4}+5-5=14-5

\frac{m}{-4}=9

-4\times \frac{m}{-4}=-4\times 9

m=-36

(H)

\frac{n}{8}-3=-11

\frac{n}{8}-3+3=-11+3

\frac{n}{8}=-8

8\times \frac{n}{8}=8\times (-8)

n=-64

(O)

\frac{w}{-5}+17=-3

\frac{w}{-5}+17-17=-3-17

\frac{w}{-5}=-20

-5\times (\frac{w}{-5})=-5\times (-20)

w=100

(T)

\frac{-x}{3}+4=20

\frac{-x}{3}+4-4=20-4

\frac{-x}{3}=16

-3\times \frac{-x}{3}=-3\times 16

48

(E)

\frac{-a}{10}-8=-24

\frac{-a}{10}-8+8=-24+8

\frac{-a}{10}=-16

-10\times \frac{-a}{10}=-10\times -16

a 160

(L)

Let the number be x. So,

\frac{x}{-7}-2=10

\frac{x}{-7}-2+2=10+2

\frac{x}{-7}=12

-7\times \frac{x}{-7}=-7\times 12

x=-84

So, the number is -84.